HETAV PATEL

Grade 11,

find intersection angle of the circle x²+y²-4x-1=0and x²+y²-2y-9=0

find intersection angle of the circle x²+y²-4x-1=0and x²+y²-2y-9=0

Grade:12

1 Answers

ASHWIN SRIKANTH
7 years ago
\cos \Theta =\left ( d^{_{2}}-r1^{_{2}}-r2^{_{2}}\left \right ) /2r1r2
here thetha is the angle between the two curves and r1 r2 are the radii of the given circles and d is the distance between their centres.now you can substitute the directly and find the answer.thank you